OpenStudy (anonymous):

Find the unknown: (-5y-2)^2+11=18

OpenStudy (bahrom7893):

take -1 out from (-5y - 2)^2:

OpenStudy (bahrom7893):

(-1[5y+2])^2 = [5y+2]^2

OpenStudy (bahrom7893):

so now: [5y+2]^2 = 18 - 11 (move 11 to the right)

OpenStudy (bahrom7893):

(5y+2)^2 = 7 25y + 20y + 4 = 7

OpenStudy (bahrom7893):

move 7 to the left: 25y + 20y +4 -7 = 0 25y + 20y - 3 = 0

OpenStudy (bahrom7893):

Now find discriminant: D = b^2 - 4ac, where b = 20, a = 25 and c = -3

OpenStudy (anonymous):

what did you do with the -1 after you factored it out?

OpenStudy (bahrom7893):

Oh it was inside a square right, so: (-1[5y+2])^2 = (-1)^2 [5y+2]^2 = 1 * [5y+2]^2 = [5y+2]^2

OpenStudy (bahrom7893):

Oh it was inside a square right, so: (-1[5y+2])^2 = (-1)^2 [5y+2]^2 = 1 * [5y+2]^2 = [5y+2]^2

OpenStudy (bahrom7893):

sorry for late response, i just keep crashing..

OpenStudy (anonymous):

no problem

OpenStudy (bahrom7893):

25y + 20y - 3 = 0 D = b^2 - 4ac, where b = 20, a = 25 and c = -3 D = 400 + 300 = 700

OpenStudy (anonymous):

i got it! thanks! :)

OpenStudy (bahrom7893):

okay np.. can you please click on become a fan, thanks!

OpenStudy (anonymous):

sure!